OpenStudy (anonymous):

To travel 147 miles, it takes Sue, riding a moped, 2 hours less time than it takes Doreen to travel 72 miles riding a bicycle. Sue travels 13 miles per hour faster than Doreen. Find the times and rates of both girls.

OpenStudy (shinalcantara):

let 'vs' be the velocity of Sue 'vd' be the velocity of Doreen \[t - 2 = \frac{147}{v_s} ----equation 1\] \[t = \frac{72}{v_d} -----equation 2\] \[v_s = v_d + 13 -----equation 3\] from equation 1 \[t - 2 = \frac{147}{v_s}\] \[t = \frac{147}{v_s} + 2 ----equation 1'\] equate equation 1' and 2 \[ \frac{147}{v_s} + 2 = \frac{72}{v_d} \] substitute equation 3 \[\frac{147}{v_d+13}+2 = \frac{72}{v_d}\] rearranging you'll have: \[2{v_d}^2 + 101v_d = 936\] \[v_d = 8mph\] substitute to equation 3 \[v_s = v_d + 13\] \[v_s = 8 + 13\] \[v_s = 21mph\]

OpenStudy (shinalcantara):

for the time: using equation 1' \[t = \frac{147}{v_s} \] \[t = \frac{147mi}{21mph}\] \[t = 7 hours ---Sue's time\] \[t = \frac{72}{v_d}\] \[t = \frac{72mi}{8mph}\] \[t = 9 hours ----Doreen's time\]